# Pure C*-algebras

### v1

@@ 5. Pure C*-algebras
If $A$ is $(m,n)$-pure, then since it has $m$-comparison Lemma 3.6 implies that $A$ has locally bounded comparison amplitude. Since it is also $n$-almost divisible, we may apply Theorem 5.7 to conclude that $A$ is pure. ∎
@@ 6. C*-algebras with the Global Glimm Property and finite nuclear dimension
The next result is our variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras. This has essentially appeared in the argument of [Vil23, Proposition 4.1].
@@
Let $A$ be a nowhere scattered $C^{*}$-algebra with $\dim_{\mathrm{nuc}}(A)\leq m$. Then for every $k\geq 1$, and every pair $x^{\prime},x\in\operatorname{Cu}(A)$ satisfying $x^{\prime}\ll x$, there exists $y\in\operatorname{Cu}(A)$ such that
@@
By [TV24, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
@@
Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24, Theorem 3.1]. It also has nuclear dimension at most $m$ (by [WZ10, Proposition 2.5]), whence we can apply [RT17, Proposition 3.2(ii)] for $k$ (and for $\overline{aAa}$, $a$ and $\varepsilon$) to obtain $b\in\overline{aAa}_{+}$ such that
@@
The following is inspired by [RT17, Lemma 3.4].
@@
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m,l\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $L=L(m,l)$ such that, for every $x^{\prime}\ll x$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
for $j=0,\ldots,l$.
@@
Let $x^{\prime},x\in\operatorname{Cu}(A)$ be such that $x^{\prime}\ll x$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
  • (i)
    $B$ has the Global Glimm Property;
  • (ii)
    $\dim_{\mathrm{nuc}}(B)\leq m$;
  • (iii)
@@
Therefore we may assume that $A$ is separable and that, as in the proof of Lemma 6.1 above, we may also assume that $A$ is stable.
Now we apply induction on $l$. Note that, if $l=0$, one can take $y_{0}:=x$. Let us assume that $l=1$.
@@
$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}.$
@@
Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2 (1)), we get
$e^{\prime}\leq 2m(2m+3)y_{1}.$
@@
$x^{\prime}\ll 2e^{\prime}\leq 4m(2m+3)y_{1}.$
Thus, setting $L:=\max\{4m+6,4m(2m+3)\}$, we see that $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$. This shows the statement for $l=1$.
Now let $l\in{\mathbb{N}}$ and assume that we have proven the statement for $l-1$.
Take $x^{\prime},x\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x$. We get $L^{\prime}\in{\mathbb{N}}$ and elements $z_{0},\ldots,z_{l-1}$ such that
$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L^{\prime}z_{j}$
@@
Set $y_{j}:=z_{j}$ for each $j\leq l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that $z_{l-1}^{\prime}\ll z_{l-1}$ and $x^{\prime}\ll L^{\prime}z_{l-1}^{\prime}$. By the argument above, one finds $L^{\prime\prime}\in{\mathbb{N}}$ and $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L^{\prime\prime}y_{l-1},L^{\prime\prime}y_{l}.$
One can now check that the elements $y_{0},\ldots,y_{l}$ and $L:=L^{\prime}L^{\prime\prime}$ satisfy the required conditions. ∎
@@
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $N=N(m)$ such that $\operatorname{Cu}(A)$ is $N$-almost divisible.
@@
The proof is inspired by that of [RT17, Lemma 3.6]. Let $L$ be the constant obtained from Lemma 6.2 for $m$ and $l=m$. We verify the statement for $N:=3L(m+1)$.
Let $x^{\prime}\ll x$ and $k\geq 1$. We need to find $y$ such that
@@
$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\leq Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
@@
and thus $k\widehat{c^{\prime}}<_{s}\widehat{y_{j}}$.
By $m$-comparison, we get
@@
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and that has nuclear dimension at most $m$. By Remarks 3.2 (1), we have that $A$ has $m$-comparison. Using Proposition 6.3, we know that $A$ is $n$-almost divisible for some $n$. Hence, $A$ is $(m,n)$-pure, and the result now follows from Theorem 5.8. ∎
Remark 6.5.
In [RT17, Theorem 3.1] it is shown that if a nowhere scattered $C^{*}$-algebra $A$ has finite nuclear dimension and no quotient of $A$ has a simple, purely infinite ideal, then $A$ is pure. Further, every approximately subhomogeneous $C^{*}$-algebra (ASH-algebra) has locally finite nuclear dimension [NW06]. By [ENST20, Theorem A], every $\mathcal{Z}$-stable ASH-algebra has decomposition rank at most two, and therefore finite nuclear dimension.
Combining these known results, we obtain another partial positive answer to Question B: For a separable, nowhere scattered ASH-algebra whose primitive ideal space has a basis of compact-open subsets or is Hausdorff, the properties ‘finite nuclear dimension’, ‘$\mathcal{Z}$-stability’ and ‘pureness’ are equivalent. If one additionally assumes that the algebra has stable rank one, then these properties are also equivalent to ‘strict comparison of positive elements’.

### v3

@@ 5. Pure C*-algebras
If $A$ is nowhere scattered and has real rank zero, then $A$ is almost divisible as noted in Paragraph 4.2. If $A$ is nowhere scattered and has stable rank one, then $A$ is functionally divisible by Proposition 4.10. Now, in both cases the result follows from Proposition 5.11. ∎
Theorem 5.13.
Let $A$ be a simple, unital, non-elementary $C^{*}$-algebra with a unique normalized quasitrace. Assume that $A$ has controlled comparison (for example, $m$-comparison for some $m$). Then $A$ is pure.
Proof.
By Proposition 4.11, $\operatorname{Cu}(A)$ is functionally divisible. Now, the result follows from Proposition 5.11. ∎
Example 5.14.
Let $G$ be an infinite, discrete group such that the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(G)$ is simple. Then $C^{*}_{\mathrm{red}}(G)$ is simple, unital, non-elementary and has a unique trace by [BKKO17, Corollary 4.3]. If $G$ is exact, then quasitraces on $C^{*}_{\mathrm{red}}(G)$ are traces by Haagerup’s theorem ([Haa14]), and it follows that $C^{*}_{\mathrm{red}}(G)$ has a unique normalized quasitrace. In this case, Theorem 5.13 applies, and it follows that $C^{*}_{\mathrm{red}}(G)$ is pure if, and only if, it has controlled comparison.
It has been proved very recently in [AGKP24] that, for $\mathbb{F}_{n}$, the free group on $n$ generators, the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(\mathbb{F}_{n})$ has strict comparison (equivalently, if it is pure). The result in fact holds for a large class of discrete groups.
For non-exact groups, we ask the following question:
Question 5.15.
Do simple reduced group $C^{*}$-algebras have a unique normalized quasitrace?
Example 5.16.
In his groundbreaking work on regularity properties for simple, nuclear $C^{*}$-algebras, Villadsen constructed examples of such algebras that fail strict comparison, and consequently are not pure. The algebras of ‘first type’ from [Vil98] have stable rank one (but possibly a complicated trace simplex), while the algebras of ‘second type’ from [Vil99] have higher stable rank but a unique tracial state (and consequently also a unique quasitracial state). Applying Propositions 4.10 and 4.11, we see that both types of Villadsen algebras are automatically functionally divisible.
We deduce that Villadsen algebras that fail strict comparison also fail the much weaker property of controlled comparison (and in particular, they do not have $m$-comparison for any $m$).
@@ 6. C*-algebras with the Global Glimm Property and finite nuclear dimension
In this section, we prove Theorem B. For this, we use that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible; see [TV23, Theorem 3.6]. We then verify a particular statement of the non-simple Toms-Winter conjecture for separable, locally subhomogeneous $C^{*}$-algebras with stable rank one and topological dimension zero, which yields Theorem E.
The next result is a variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras. This has essentially appeared in the argument of [Vil23, Proposition 4.1]. We use $\dim_{\mathrm{nuc}}(A)$ to denote the nuclear dimension of $A$.
@@
Let $m\in{\mathbb{N}}$, and let $A$ be a nowhere scattered $C^{*}$-algebra satisfying $\dim_{\mathrm{nuc}}(A)\leq m$. Then for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in\operatorname{Cu}(A)$ such that
@@
By [TV24a, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
@@
Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24a, Theorem 3.1]. It also has nuclear dimension at most $m$ (by [WZ10, Proposition 2.5]), whence we can apply [RT17, Proposition 3.2(ii)] for $k$ (and for $\overline{aAa}$, $a$ and $\varepsilon$) to obtain $b\in\overline{aAa}_{+}$ such that
@@
The next two results are inspired by [RT17, Lemma 3.4].
@@
Let $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra with the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$. Set $L:=14m^{2}+6$. Then for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, there exist $y_{0},y_{1}\in\operatorname{Cu}(A)$ such that
$y_{0}+y_{1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{0},Ly_{1}.$
@@
Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
  1. (i)
    $B$ has the Global Glimm Property and $\dim_{\mathrm{nuc}}(B)\leq m$;
  2. (ii)
@@
Identifying $\operatorname{Cu}(B)$ with its image in $\operatorname{Cu}(A)$, we view $x^{\prime},x$ as elements in $\operatorname{Cu}(B)$, and it suffices to find $y_{0},y_{1}\in\operatorname{Cu}(B)$ with $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$. Therefore we may assume that $A$ is separable.
@@
$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}\leq(14m^{2}+6)y_{0}=Ly_{0}.$
@@
Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2), we get
$e^{\prime}\leq(m+1)(2m+3)y_{1}.$
@@
$x^{\prime}\ll 2e^{\prime}\leq 2(m+1)(2m+3)y_{1}=2(2m^{2}+5m+3)y_{1}\leq(14m^{2}+6)y_{1}=Ly_{1}.$
This shows that $y_{0}$ and $y_{1}$ have the desired properties. ∎
Lemma 6.3.
Given $m,l\in{\mathbb{N}}$, set $L:=(14m^{2}+6)^{l}$. Then, for every $C^{*}$-algebra $A$ with the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$, and for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
for $j=0,\ldots,l$.
Proof.
Fix some $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra that has the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$. Write $L(m,l):=(14m^{2}+6)^{l}$ for $l\geq 0$.
By induction over $l$, we verify that the following statement holds:
  1. ($D_{l}$)
    For every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
    $y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l)y_{j}$
    for $j=0,\ldots,l$.
To verify ($D_{0}$), given $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ we simply use $y_{0}=x$. Further, ($D_{1}$) was shown in Lemma 6.2. Now let $l\in{\mathbb{N}}$ with $l\geq 2$, and assume that we have proved ($D_{l-1}$). To verify ($D_{l}$), let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$. Applying ($D_{l-1}$) for $x^{\prime}\ll x$, we get $z_{0},\ldots,z_{l-1}\in\operatorname{Cu}(A)$ such that
$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l-1)z_{j}$
@@
Set $y_{j}:=z_{j}$ for $j=0,\ldots,l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that
$z_{l-1}^{\prime}\ll z_{l-1},\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l-1)z_{l-1}^{\prime}.$
Applying ($D_{1}$) for $z_{l-1}^{\prime}\ll z_{l-1}$, we find $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L(m,1)y_{l-1},L(m,1)y_{l}.$
Then $y_{0}+\ldots+y_{l}\leq x$, and
$x^{\prime}\ll L(m,l-1)z_{j}\leq L(m,l)z_{j}=L(m,l)y_{j}$
for $j=0,\ldots,l-2$ and
$x^{\prime}\ll L(m,l-1)z_{l-1}^{\prime}\ll L(m,l-1)L(m,1)y_{j}=L(m,l)y_{j}$
for $j=l-1,l$. This shows that $y_{0},\ldots,y_{l}$ have the desired properties. ∎
@@
Given $m\in{\mathbb{N}}$, there exists $N=N(m)$ with the following property: For every $C^{*}$-algebra $A$ with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$, the Cuntz semigroup $\operatorname{Cu}(A)$ is $N$-almost divisible.
@@
The proof is inspired by that of [RT17, Lemma 3.6]. Let $L=L(m,m)$ be the constant obtained from Lemma 6.3. We verify the statement for $N:=3L(m+1)$.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$. To show that $\operatorname{Cu}(A)$ is $N$-almost divisible, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $k\geq 1$. We need to find $y\in\operatorname{Cu}(A)$ such that
@@
$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
@@
and thus $kc^{\prime}<_{s}y_{j}$.
As noted in Remarks 3.2, $A$ has $m$-comparison, and we get
@@
Let $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$. Then $A$ has $m$-comparison, as noted in Remarks 3.2. Further, by Proposition 6.4, we know that $A$ is $n$-almost divisible for some $n$. Hence, $A$ is $(m,n)$-pure, and the result now follows from Theorem 5.7. ∎
Theorem 6.6.
Let $A$ be a separable, locally subhomogeneous $C^{*}$-algebra that has stable rank one and topological dimension zero. Then the following statements are equivalent:
  1. (1)
    $A$ has the Global Glimm Property and finite nuclear dimension.
  2. (2)
    $A$ is $\mathcal{Z}$-stable.
  3. (3a)
    $A$ is pure.
  4. (3b)
    $A$ is nowhere scattered and has strict comparison of positive elements.
Proof.
By Theorem 6.5, (1) implies (3a). By Proposition 5.2, (2) implies (3a), and (3a) implies (3b). By Theorem 5.12, (3b) implies (3a).
Every locally subhomogeneous $C^{*}$-algebra has locally finite nuclear dimension by [NW06]. Since locally subhomogeneous $C^{*}$-algebras have no simple, purely infinite ideal-quotients, it follows from [RT17, Theorem 7.10] that (3a) implies (2).
Finally, by [ENST20, Theorem A], every $\mathcal{Z}$-stable, locally subhomogeneous $C^{*}$-algebra has decomposition rank at most two, and therefore finite nuclear dimension. Further, by Proposition 5.2, $\mathcal{Z}$-stability implies the Global Glimm Property. This shows that (2) implies (1). ∎